### 04 Mar 2026 v2

@@ Abstract
Regular black holes are often used to address singularities, but they typically involve a potentially unstable de Sitter core and an internal horizon that breaks predictability. Integrable singularities (IS) have recently gained attention because they avoid both issues and exhibit finite tidal forces, allowing nondestructive radial infall. First, we present a new BH solution sourced by a string fluid (FS) that exhibits an IS. Motivated by the divergence of the conserved energy in the cloud of strings (CS) model, we introduce an energy density profile based on the screening of the CS energy density within an FS framework, thereby yielding finite conserved energy. On the other hand, the idea that an interior region, rather than a pointlike mass, can generate a Schwarzschild black hole exterior region has recently gained attention Ovalle (2024); Maeda (2025). This is obtained by matching an interior solution to the Schwarzschild exterior at the event horizon. Secondly, motivated by the variety of singular BH solutions in the literature, we establish the conditions that an interior region with an IS must satisfy in order to represent the interior of a generic exterior BH solution, with Schwarzschild being only a particular case of the latter. We derive the junction conditions (JC) between the interior and exterior regions, showing that they lead to temperature continuity at the interface, while discontinuities in tangential pressure lead to phase transitions. We propose that the nature of the interior region is described by CS and FS, while the exterior corresponds to Reissner–Nordström.

### 05 Mar 2026 v3

@@ Abstract
BHs with integrable singularities (IS) have gained attention because, unlike usual regular BHs, they avoid a potentially unstable de Sitter core and the presence of an internal horizon that breaks predictability, while exhibiting finite tidal forces that allow nondestructive radial infall. First, we present a new BH solution sourced by a string fluid (FS) exhibiting an IS. We introduce an energy density profile obtained by screening the cloud of strings (CS) density within an FS framework, leading to finite conserved energy, unlike the CS case where it diverges. On the other hand, the idea that an interior region, rather than a pointlike mass, can generate a Schwarzschild exterior BH has recently gained attention Ovalle (2024); Maeda (2025). This is achieved by matching an interior region to the Schwarzschild exterior at the event horizon. Motivated by the variety of singular BH solutions in the literature, we establish the conditions that an interior region with an IS must satisfy to represent the interior of a generic exterior BH solution, with Schwarzschild as only a particular case. We derive the junction conditions between the interior and exterior regions, showing that they ensure temperature continuity at the interface, while discontinuities in tangential pressure signal phase transitions. We examine an interior described by CS and FS, matched to a Reissner–Nordström exterior.